OpenStudy (anonymous):

what is n and x if √(x/10)=√(2n-6)? (can somebody help me???)

OpenStudy (ash2326):

we have \[\sqrt (x/10)=\sqrt ( 2n-6)\] fora sqare root , we know that 2n-6>=0 so n>=3 if n=3 then x=0 let's find other solutions let's put any n>3 say n=5 then root(x/10)=root(2n-6) or x/10=2n-6 x=10(2*5-6) x=40 similarly we can find many values for n and x, but n>=3